Help please !
ime Remaining: 01 40 39 Submit Te This Question: 1 pt 34 of 34 (13 complete) This Test: 40 pts possi The value given below is discrete. Use the continuity correction and describe the region of the normal dst Probability of fewer than 7 passengers who do not show up for a flight Choose the correct answer below O A. The area to the right of 7.5 O B. The area to the left of 6.5 O C. The area to the right of 6.5 O D. The area to the left of 7.5 O E. The area between 6.5 and 7.5 Click to select your answer l Permissions 1 con
Answer #1

Here we need to find the probability P(X < 7).

Using continuity correction factor, the probability we need to find is

P(X < 7) = P(X < 7-0.5) = P(X < 6.5)

Answer: Correct option is B.